OpenStudy (anonymous):

find the monic quadratic polynomial that crosses the y axis at (0,-6) and the x axis at (3,0)

OpenStudy (anonymous):

You know that 3 is a root so can you just do (x - 3)(x - a) = x^2 - ax - 3x + 3a Then when x = 0, y = -6 3a = - 6 a = -2 so x^2 -x -6

OpenStudy (anonymous):

THANK YOU VERY MUCH!

OpenStudy (anonymous):

umm how does finding a=-2 help you deduce x^2 -x -6

OpenStudy (anonymous):

oh dw i get it now haha

OpenStudy (anonymous):

Sub the values into x^2 - ax - 3x + 3a x^2 + 2x - 3x +3(-2) = x^2 -x -6

OpenStudy (anonymous):

Oh ok. :)